a 3M BAR is used to lift a stone. A 90 kg person exerts all his weight to one end of the bar which is 2 1/2 m from the fulcrum and balances the stone. The weight of the stone is approximately?
PLEASE I AM ASKING FOR HELP, WHO CAN ANSWER MY PRAYER.......

Answer by stanbon(48535)

a 3M BAR is used to lift a stone. A 90 kg person exerts all his weight to one end of the bar which is 2 1/2 m from the fulcrum and balances the stone. The weight of the stone is approximately?
------------------------
weight*length = weight*length
90*(5/2) = x*3
30(5/2) = x
x = 75 kg (the weight of the stone)
=====================================================
Cheers,
Stan H.
=====================================================